find equation of a line Through:(3,5)with slope=5/3
Answer by lawdej(58)

y-y1=m(x-x1), m=slpoe
(x1,y1)=(3,5),m=5/3
y-5=5/3(x-3)
3y-15=5x-15
3y=5x-15+15
3y=5x
y=5/3 x